# Alchemy: Amplifying Theorem-Proving Capability through Symbolic Mutation

### v1

@@ Abstract
Formal proofs are challenging to write even for experienced experts. Recent progress in Neural Theorem Proving (NTP) shows promise in expediting this process. However, the formal corpora available on the Internet are limited compared to the general text, posing a significant data scarcity challenge for NTP. To address this issue, this work proposes Alchemy, a general framework for data synthesis that constructs formal theorems through symbolic mutation. Specifically, for each candidate theorem in Mathlib, we identify all invocable theorems that can be used to rewrite or apply to it. Subsequently, we mutate the candidate theorem by replacing the corresponding term in the statement with its equivalent form or antecedent. As a result, our method increases the number of theorems in Mathlib by an order of magnitude, from 110k to 6M. Furthermore, we perform continual pretraining and supervised finetuning on this augmented corpus for large language models. Experimental results demonstrate the effectiveness of our approach, achieving a 5% absolute performance improvement on Leandojo benchmark. Additionally, our synthetic data achieve a 2.5% absolute performance gain on the out-of-distribution miniF2F benchmark. To provide further insights, we conduct a comprehensive analysis of synthetic data composition and the training paradigm, offering valuable guidance for developing a strong theorem prover.
@@ 1 Introduction
The key challenge of theorem-proving lies in data scarcity (Li et al., 2024; Trinh et al., 2024). Due to the difficulties associated with the manual formalization of theorems, formal corpora available on the Internet are relatively scarce compared to the general text (Azerbayev et al., 2023). Synthetic data has shown promise in alleviating the data scarcity problem. Some works propose to directly create theorems in symbolic space. For instance, Wang & Deng (2020) attempts to train a neural theorem generator on human-written formal theorems for the low-weighted formal system Metamath. Other efforts focus on generating theorems based on symbolic rules (Wu et al., 2021; Trinh et al., 2024), which are restricted to a specific domain of mathematics, such as inequality theorems and 2D geometry. Additionally, there are endeavors focused on autoformalization (Xin et al., 2024; Ying et al., 2024), which typically translates natural language mathematical problems into formalized statements, samples correct proofs, and retrains the theorem prover iteratively. Autoformalization has yielded promising results in competition-level theorem-proving tasks through the use of large autoformalized datasets. However, the process of formalizing problems and retrieving proofs is labor-intensive and cost-prohibitive. The distribution of formalized theorems is constrained by the pool of human-collected natural language problems and the intrinsic capabilities of the model. Compared to autoformalization, synthesizing theorems in symbolic space is a more direct process without intermediate translation, and is also easier to scale up to large, cost-effective CPU units.
@@
Building upon the advanced Lean theorem prover, we introduce a general method that synthesizes theorems directly in symbolic space. We analogize theorem synthesis to constructing functions in general programming language and adopt an up-to-down approach. Initially, a new statement (function declaration) is constructed for each candidate theorem. Specifically, with the mathematical library of Lean Mathlib4111https://github.com/leanprover-community/mathlib4 as seed data, we aim to find a symbolic manipulation $\Phi$ between two existing statements. We posit that Lean’s tactics serve as suitable candidates for manipulation $\Phi$ because of their efficacy in handling symbolic expressions. $\{rw,apply\}$ are basic tactics frequently used in theorem proving and capable of handling the equality and implication relationship between terms. We assign both tactics to the set of manipulations $\Phi$ and retrieve the invocable theorems for each candidate theorem by executing a predefined list of instructions in an interactive Lean environment. Then we mutate the candidate statement by replacing its components with their corresponding equivalent forms or logical antecedents. Ultimately, we construct the corresponding proof (function body) based on the existing proof and verify its correctness using Lean. The worked example shown in Fig 1 illustrates the entire procedure of our algorithm. This algorithm is executed on a large CPU-only computing unit for several days. Our method increases the number of theorems in Mathlib by an order of magnitude from 110,657 to 6,326,679. This significant increase in the number of theorems demonstrates the potential of creating theorems in symbolic space.
@@
We pre-train the LLMs on the combination of Mathlib theorems and their mutated variants. Then we fine-tune the models on the extracted state-tactic pairs, composing both the training split of Mathlib and additional synthesized state-tactic pairs. We demonstrate the effectiveness of our method by evaluating the theorem-proving capability of these provers on the challenging Leandojo benchmark. Our synthetic data improve the performance by around 5% (over 70 theorems) on the novel_premises split. Furthermore, the synthesized data exhibit promise in enhancing the out-of-distribution theorem-proving ability of LLMs, as evidenced by a performance increase of about 2.5% on the competition-level miniF2F benchmark.
@@
Our main contributions are as follows. 1) To the best of our knowledge, this work represents the first general data synthesis framework in the symbolic space for the Lean theorem prover, effectively complementing mainstream autoformalization-based methods. Notably, our synthesis pipeline increases the number of theorems in Mathlib4 by an order of magnitude. 2) The synthesized data and associated code will be made open-source to facilitate further research in data synthesis for formal systems. Also, the synthesized theorems can serve as a valuable supplement to Mathlib. 3) We conduct a comprehensive evaluation on both in-distribution and out-of-distribution benchmarks, providing empirical insights to enhance the theorem-proving capabilities of LLMs.
@@ 2 Related Work
Proof assistants such as Lean (de Moura et al., 2015), Isabelle (Paulson, 1994) or Coq (Barras et al., 1997) are gaining traction within the mathematical community. These tools help mathematicians in interactively formalizing and checking the correctness of proofs (Tao, 2024). Neural networks have shown promise in lowering the barrier of using a specific formal language for mathematicians, serving as a copilot (Song et al., 2024; Welleck & Saha, 2023). Polu & Sutskever (2020) propose to prove theorems automatically by training a decoder-only transformer to predict the next proofstep and construct the entire proof through a predefined search tragedy. Then a series of works seek to enhance the efficiency of this framework by incorporating auxiliary training objectives (Han et al., 2022), conducting reinforcement learning (Polu et al., 2023; Xin et al., 2024), improving proof search tragedy (Lample et al., 2022; Wang et al., 2023; Xin et al., 2024), refining the premise-selection (Mikula et al., 2023; Yang et al., 2023) and so on. Our work follows the framework proposed by Polu & Sutskever (2020), using proofstep prediction as the objective and best-first-search as the search tragedy.
Synthetic Theorem Creation
@@
Data scarcity is a main challenge for NTP (Li et al., 2024). Synthetic data can effectively alleviate this problem alongside manual data collection (Wu et al., 2024). The current approach for synthesizing theorems diverges into two pathways. For autoformalization-based methods, the prevalent statement-level autoformalization is to translate a set of natural language problems into formal statements, followed by expert iteration to sample a collection of proofs for these statements (Wu et al., 2022; Xin et al., 2024; Ying et al., 2024). The proof-level autoformalization (Jiang et al., 2023; Huang et al., 2024) leverages LLM to generate a proof sketch, which is completed by symbolic engines such as Sledgehammer (Böhme & Nipkow, 2010). In contrast, the second pathway focuses on synthesizing theorems in formal space. Wang & Deng (2020) propose to train a neural theorem generator to synthesize theorems on a low-weight formal system, Metamath (Megill & Wheeler, 2019) which has only one tactic substitute. Wu et al. (2021) sequentially edits the seed expression according to a predefined set of axioms and an axiom order to create a new statement, concatenating the implications from all steps to build a complete proof. This method is used to create theorems on domains grounded in well-established axioms, such as inequality theorems and ring algebra (Polu & Sutskever, 2020). Beyond these works, AlphaGeometry (Trinh et al., 2024) can solve olympiad geometry without human demonstrations by constructing statements and proofs in symbolic space from scratch, using a carefully designed deduction engine and large-scale computing resources. Our method aims to directly synthesize theorems in symbolic space on the advanced Lean theorem prover, fully utilizing the power of computing.
Benchmarks for Theorem Proving
@@
Most neural theorem provers based on Lean are primarily trained on Lean’s mathematical library, Mathlib. It encompasses a broad spectrum of mathematical subjects (e.g., algebra and analysis), composed of over 120,000 theorems along with their respective axioms and definitions. Researchers test the capability of neural models to prove in-distribution theorems on a held-out set of Mathlib (Polu & Sutskever, 2020; Han et al., 2022; Polu et al., 2023). Yang et al. (2023) creates a challenging data split of Mathlib (novel_premise split) which requires testing proofs to use at least one premises not seen in the training stage and mitigates the over-estimated phenomena in the traditional setting of evaluation (random split). Another widely-used benchmark, miniF2F, (Zheng et al., 2022) is a cross-system benchmark and includes competition-level problems as well as IMO-level problems in the domain of algebra and number theory.
@@ 3 Method
Theorems written in Lean can be viewed as a special form of code, where declarations and function bodies possess precise mathematical meanings. The initial step in creating a new theorem involves formulating a theorem statement (function declaration) that defines the essence of the theorem. Then, one must verify its correctness by generating a proof block (function body) and submitting it to the proof assistant for validation. The resulting theorems that pass type checking can serve as supplementary data for training a neural theorem prover.
@@ 3.1 Statement Generation
Constructing a new statement is the first step in creating a Lean theorem. The candidate theorem $t$ has a statement denoted as $s$. In the corresponding Lean repository, there exists a set of potentially invocable theorems $T_{p}=\{t_{i}\}_{i=0}^{N}$. We assume that the challenge in creating a new theorem involves effectively leveraging the possibly invocable theorem $t_{i}$ to mutate the candidate statement $s$. This understanding arises from two perspectives. Each theorem in Lean can be represented in the form of a proof tree as presented in Fig 1. The leaf nodes represent the assumptions, and the root node signifies the assertion. At the tree level, the task of generating a new Lean theorem with existing theorems is equivalent to defining operations $\Phi$ that combine the proof trees of $t_{i}$ and $t$. To streamline this process, our focus is solely on establishing the connection between the root node of $t_{i}$ and the leaf node (or root node) of the candidate theorem $t$. From a mathematical standpoint, we can transform a target formula into an equal variant or break it down into multiple subformulas that suffice to prove the original formula, by employing the equality or “only if” relationship between formulas. The mathematical interconnections between formulas provide heuristic insights on how to mutate $s$ to create a new theorem. Similarly, we can substitute the terms in $s$ with their equivalent forms or logical antecedents. For instance, consider the statement $a+b>c+d,m>0\rightarrow m(a+b)>m(c+d)$ and the known theorems $a>b\iff e^{a}>e^{b}$ and $a>c,b>d\implies a+b>c+d$. From these, we can derive new theorems: $a+b>c+d,m>0\rightarrow e^{m(a+b)}>e^{m(c+d)}$, and $a>c,b>d,m>0\implies m(a+b)>m(c+d)$. In summary, identifying operations $\Phi$ that use $t_{i}$ to modify the assumptions or assertion of $s$ is the primary step in constructing new statements.
@@
With their intrinsic mathematical meanings and proficiency in manipulating terms within Lean, tactics are promising candidates for the operations $\Phi$. Following the preceding discussion, we choose two frequently used basic tactics, rw and apply to formulate $\Phi$.
@@
  • rw The “rewriting” tactic rw is mostly used to replace some terms in the target expression with their equivalent forms according to the given identity or iff (a.k.a., if and only if) rules [^2] Strictly speaking, the rw tactic is used to handling equality in Lean, the identity and iff are just some kinds of equality.. In the presence of an identity $h:a=b$ or an iff rule $h:P\iff Q$, rw [h] substitutes all occurrences of term on the left side of equality in the proof goal with term on the right side. The direction of substitution can be reversed by adding a back arrow in the bracket (rw [$\leftarrow$ h]). The target of rewriting can also be changed using at, e.g. rw [h] at $h_{1}$, where $h_{1}$ is an arbitrary assumption of the current proof state.
@@
  • apply The apply tactic is a “suffice-to” tactic. Given an implication, it will match the consequent with the proof goal. If matched, it will transform the goal into the antecedent of the implication. With an implication rule $h:P\implies Q$ and a proof goal $Q$, then apply [h] will reduce the goal to proving $P$, which means that “proving P suffices to prove Q by implication”. Similarly, apply can be used to modify the assumption by deducing the implication forward. With assumption $h_{1}:P$, then apply [h] at $h_{1}$ will change $h_{1}$ into $Q$, which means “If P is true, then we can assert Q is true by the implication”.
@@
Input: candidate statement $s$, potential invocable theorems $T_{p}$, instruction templates $I$
Output: invocable theorems $T_{i}$triangleright $T_{i}:\{(init\_state,next\_state,instruction)\cdots\}$
@@
              Add $(init\_state,next\_state,inst)$ to $T_{i}$
@@
To generate a new statement, we need to find the relationship between the candidate statement $s$ and the potentially invocable theorems $T_{p}$. The pseudocode outlined in Algorithm 1 describes the main procedure to find invocable theorems. The process involves initializing a gym-like environment to interact with Lean and extracting the initial proof state for the candidate statement. Then, the algorithm iteratively tests whether one theorem can be used to rewrite or apply to the candidate theorem leveraging the instruction templates shown in Table 1. Suppose the feedback from the interactive environment is deemed valid according to predefined criteria, the algorithm adds the proof states before and after the tactic running together with the respective instruction to the set of invocable theorems $T_{i}$. More information about this process is described in Appendix C.2.
@@
rw Equality invocable_theorem : $\mathbf{a=b}$ or $\mathbf{a\iff b}$
@@
apply Implication invocable_theorem : $\mathbf{a\implies b}$
@@
Mutate statements
After obtaining the initial set of invocable theorems, we applied some filtering rules to $T_{i}$ to improve the quality of the data and lower the complexity of mutating statements. With filtered invocable theorems $T_{i}$, we construct new statements by replacing the components with their equivalent forms or antecedents. Since we use tactics in Lean to formulate the operations $\Phi$, most symbolic manipulations are bypassed to the Lean proof assistant. What remains is just parsing and replacing. Specifically, for the candidate statement $s$ and instruction $i$, we utilize its abstract syntax tree to pinpoint the exact location within the code that requires modification. Then we replace the corresponding parts with mutants parsing from the subsequent proof state generated by the execution of a specific tactic. The details of our algorithm and helpful source code are described in C.3.

### v2

@@ Abstract
Formal proofs are challenging to write even for experienced experts. Recent progress in Neural Theorem Proving (NTP) shows promise in expediting this process. However, the formal corpora available on the Internet are limited compared to the general text, posing a significant data scarcity challenge for NTP. To address this issue, this work proposes Alchemy, a general framework for data synthesis that constructs formal theorems through symbolic mutation. Specifically, for each candidate theorem in Mathlib, we identify all invocable theorems that can be used to rewrite or apply to it. Subsequently, we mutate the candidate theorem by replacing the corresponding term in the statement with its equivalent form or antecedent. As a result, our method increases the number of theorems in Mathlib by an order of magnitude, from 110k to 6M. Furthermore, we perform continual pretraining and supervised finetuning on this augmented corpus for large language models. Experimental results demonstrate the effectiveness of our approach, achieving a 4.70% absolute performance improvement on Leandojo benchmark. Additionally, our approach achieves a 2.47% absolute performance gain on the out-of-distribution miniF2F benchmark based on the synthetic data. To provide further insights, we conduct a comprehensive analysis of synthetic data composition and the training paradigm, offering valuable guidance for developing a strong theorem prover. 111  The code is available at https://github.com/wclsn/Alchemy.
@@ 1 Introduction
The key challenge of theorem-proving lies in data scarcity (Li et al., 2024; Trinh et al., 2024). Due to the difficulties associated with the manual formalization of theorems, formal corpora available on the Internet are relatively scarce compared to the general text (Azerbayev et al., 2024). Synthetic data has shown promise in alleviating the data scarcity problem. Some works propose to directly create theorems in symbolic space. For instance, Wang & Deng (2020) attempts to train a neural theorem generator on human-written formal theorems for the low-weighted formal system Metamath. Other efforts focus on generating theorems based on symbolic rules (Wu et al., 2021; Trinh et al., 2024), which are restricted to a specific domain of mathematics, such as inequality theorems and 2D geometry. Additionally, there are endeavors focusing on autoformalization (Xin et al., 2024; Ying et al., 2024), which typically translates natural language mathematical problems into formalized statements, samples correct proofs, and retrains the theorem prover iteratively. Autoformalization has yielded promising results in competition-level theorem-proving tasks through the use of large autoformalized datasets (Xin et al., 2024). However, the process of formalizing problems and retrieving proofs is labor-intensive and cost-prohibitive. The distribution of formalized theorems is constrained by the pool of human-collected natural language problems and the intrinsic capabilities of the model. Compared to autoformalization, synthesizing theorems in symbolic space is a more direct process without intermediate translation, and is also easier to scale up to large, cost-effective CPU units.
@@
Building upon the advanced Lean theorem prover (de Moura et al., 2015), we introduce a general method that synthesizes theorems directly in symbolic space. We analogize theorem synthesis to constructing functions in general programming language and adopt an up-to-down approach. Initially, a new statement (function declaration) is constructed for each candidate theorem. Specifically, with the mathematical library of Lean Mathlib [^2] https://github.com/leanprover-community/mathlib4 as seed data, we aim to find a symbolic manipulation between two existing statements. We posit that Lean’s tactics serve as suitable candidates for manipulation because of their efficacy in handling symbolic expressions. $\{rw,apply\}$ are two basic tactics frequently used in theorem proving and capable of handling the equality and implication relationship between terms. We assign both tactics to the set of manipulations and retrieve the invocable theorems for each candidate theorem by executing a predefined list of instructions in an interactive Lean environment. Then we mutate the candidate statement by replacing its components with their corresponding equivalent forms or logical antecedents. Ultimately, we construct the corresponding proof (function body) based on the existing proof and verify its correctness using Lean. The worked example shown in Fig.1 illustrates the entire procedure of our algorithm. This algorithm is executed on a large CPU-only computing unit for several days. Our method increases the number of theorems in Mathlib by an order of magnitude from 110,657 to 6,326,649. This significant increase in the number of theorems demonstrates the potential of creating theorems in symbolic space.
@@
We pre-train the LLMs on the combination of Mathlib theorems and their mutated variants. Then we fine-tune the models on the extracted state-tactic pairs, composing both the training split of Mathlib and additional synthesized state-tactic pairs. We demonstrate the effectiveness of our method by evaluating the theorem-proving capability of these provers on the challenging Leandojo benchmark (Yang et al., 2023). Our synthetic data improve the performance by 4.70% (over 70 theorems) on the novel_premises split. Furthermore, the synthesized data exhibit promise in enhancing the out-of-distribution theorem-proving ability of LLMs, as evidenced by a performance increase of about 2.47% on the competition-level miniF2F benchmark (Zheng et al., 2022).
@@
Our main contributions are as follows. To the best of our knowledge, this work represents the first general data synthesis framework in the symbolic space for the Lean theorem prover, effectively complementing mainstream autoformalization-based methods. Notably, our synthesis pipeline increases the number of theorems in Mathlib by an order of magnitude. Associated code has been made open-source to facilitate further research in data synthesis for formal systems. Also, the synthesized theorems can serve as a valuable supplement to Mathlib. We conduct a comprehensive evaluation on both in-distribution and out-of-distribution benchmarks, providing empirical insights to enhance the theorem-proving capabilities of LLMs.
@@ 2 Related Work
Neural Theorem Proving. Proof assistants such as Lean (de Moura et al., 2015), Isabelle (Paulson, 1994) or Coq (Barras et al., 1997) are gaining traction within the mathematical community. These tools help mathematicians in interactively formalizing and checking the correctness of proofs (Tao, 2024). Neural networks have shown promise in lowering the barrier of using a specific formal language for mathematicians, serving as a copilot (Song et al., 2024; Welleck & Saha, 2023). Polu & Sutskever (2020) propose to prove theorems automatically by training a decoder-only transformer to predict the next proofstep and construct the entire proof through a predefined search tragedy. Then a series of works seek to enhance the efficiency of this framework by incorporating auxiliary training objectives (Han et al., 2022), conducting reinforcement learning (Polu et al., 2023; Xin et al., 2024), improving proof search tragedy (Lample et al., 2022; Wang et al., 2023; Xin et al., 2024), refining the premise-selection (Mikula et al., 2024; Yang et al., 2023) and so on.
@@
Synthetic Theorem Creation. Data scarcity is a main challenge for NTP (Li et al., 2024). Synthetic data can effectively alleviate this problem alongside manual data collection (Wu et al., 2024). The current approach for synthesizing theorems diverges into two pathways. For autoformalization-based methods, the prevalent statement-level autoformalization is to translate a set of natural language problems into formal statements, followed by expert iteration to sample a collection of proofs for these statements (Wu et al., 2022; Xin et al., 2024; Ying et al., 2024). The proof-level autoformalization (Jiang et al., 2023; Huang et al., 2024) leverages LLM to generate a proof sketch, which is completed by symbolic engines such as Sledgehammer (Böhme & Nipkow, 2010). In contrast, the second pathway focuses on synthesizing theorems in formal space. Wang & Deng (2020) propose to train a neural theorem generator to synthesize theorems on a low-weight formal system, Metamath (Megill & Wheeler, 2019) which has only one tactic substitute. Wu et al. (2021) sequentially edits the seed expression according to a predefined set of axioms and an axiom order to create a new statement, concatenating the implications from all steps to build a complete proof. This method is used to create theorems on domains grounded in well-established axioms, such as inequality theorems and ring algebra (Polu & Sutskever, 2020). Beyond these works, AlphaGeometry (Trinh et al., 2024) can solve olympiad geometry without human demonstrations by constructing statements and proofs in symbolic space from scratch, using a carefully designed deduction engine and large-scale computing resources. Our method aims to directly synthesize theorems in symbolic space on the advanced Lean theorem prover, fully utilizing the power of computing.
@@
Benchmarks for Theorem Proving. Most neural theorem provers based on Lean are primarily trained on Lean’s mathematical library, Mathlib. It encompasses a broad spectrum of mathematical subjects (e.g., algebra and analysis), composed of over 110,000 theorems along with their respective axioms and definitions. Researchers test the capability of neural models to prove in-distribution theorems on a held-out set of Mathlib (Polu & Sutskever, 2020; Han et al., 2022; Polu et al., 2023). Yang et al. (2023) creates a challenging data split of Mathlib (novel_premise split) which requires testing proofs to use at least one premises not seen in the training stage and mitigates the over-estimated phenomena in the traditional setting of evaluation (random split). Another widely-used benchmark, miniF2F, (Zheng et al., 2022) is a cross-system benchmark and includes competition-level problems as well as IMO-level problems in the domain of algebra and number theory.
@@ 3 Method
Theorems written in Lean can be viewed as a special form of code, where declarations and function bodies possess precise mathematical meanings. The initial step in creating a new theorem involves formulating a theorem statement (function declaration) that defines the essence of the theorem. Then, one must verify its correctness by generating a proof block (function body) and submitting it to the proof assistant for validation. The resulting theorems that pass type checking can serve as supplementary data for training a neural theorem prover. Following Polu & Sutskever (2020), we use proofstep prediction as the training objective and best-first-search as the search tragedy.
@@ 3.1 Statement Generation
Find invocable theorems. Constructing a new statement is the first step in creating a Lean theorem. The candidate theorem $t$ has a statement denoted as $s$. In the corresponding Lean repository, there exists $M$ potentially invocable theorems $T_{pinv}=\{t_{j}\}_{j=0}^{M-1}$. We assume that the challenge in creating a new theorem involves effectively leveraging the possibly invocable theorem $t_{j}$ to mutate the candidate statement $s$. This understanding arises from two perspectives. Each theorem in Lean can be represented in the form of a proof tree as presented in Fig.1. The leaf nodes represent the assumptions, and the root node signifies the assertion. At the tree level, the task of generating a new Lean theorem with existing theorems is equivalent to defining manipulations $\Phi$ that combine the proof trees of $t_{j}$ and $t$. To streamline this process, our focus is solely on establishing the connection between the root node of $t_{j}$ and the leaf node (or root node) of the candidate theorem $t$. From a mathematical standpoint, we can transform a target formula into an equal variant or break it down into multiple subformulas that suffice to prove the original formula, by employing the equality or “only if” relationship between formulas. The mathematical interconnections between formulas provide heuristic insights on how to mutate $s$ to create a new theorem. Similarly, we can substitute the terms in $s$ with their equivalent forms or logical antecedents. For instance, consider the statement $a+b>c+d,m>0\rightarrow m(a+b)>m(c+d)$ and the known theorems $a>b\iff e^{a}>e^{b}$ and $a>c,b>d\implies a+b>c+d$. From these, we can derive new theorems: $a+b>c+d,m>0\rightarrow e^{m(a+b)}>e^{m(c+d)}$, and $a>c,b>d,m>0\implies m(a+b)>m(c+d)$. In summary, identifying manipulations $\Phi$ that use $t_{j}$ to modify the assumptions or assertion of $s$ is the primary step in constructing new statements.
@@
With their intrinsic mathematical meanings and proficiency in manipulating terms within Lean, tactics are promising candidates for the manipulations $\Phi$. Following the preceding discussion, we choose two frequently used basic tactics, rw and apply to formulate $\Phi$.
@@
  • rw. The “rewriting” tactic rw is mostly used to replace some terms in the target expression with their equivalent forms according to the given identity or iff (a.k.a., if and only if) rules [^3] Strictly speaking, the rw tactic is used to handle equality in Lean. The identity and iff are just some kinds of equality.. In the presence of an identity $h:a=b$ or an iff rule $h:P\iff Q$, rw [h] substitutes all occurrences of term on the left side of equality in the proof goal with term on the right side. The direction of substitution can be reversed by adding a back arrow in the bracket (rw [$\leftarrow$ h]). The target of rewriting can also be changed using at, e.g. rw [h] at $h_{1}$, where $h_{1}$ is an arbitrary assumption of the current proof state.
@@
  • apply. The apply tactic is a “suffice-to” tactic. Given an implication, it will match the consequent with the proof goal. If matched, it will transform the goal into the antecedent of the implication. With an implication rule $h:P\implies Q$ and a proof goal $Q$, then apply [h] will reduce the goal to proving $P$, which means that “proving P suffices to prove Q by implication”. Similarly, apply can be used to modify the assumption by deducing the implication forward. With assumption $h_{1}:P$, then apply [h] at $h_{1}$ will change $h_{1}$ into $Q$, which means “If P is true, then we can assert Q is true by the implication”.
@@
Input: candidate statement $s$, potential invocable theorems $T_{pinv}$, instruction templates $I$
Output: invocable theorems $T_{inv}$triangleright $T_{inv}:\{(init\_state,next\_state,instruction)\cdots\}$
@@
              Add $(init\_state,next\_state,inst)$ to $T_{inv}$
@@
To generate a new statement, we need to find the relationship between the candidate statement $s$ and the potentially invocable theorems $T_{pinv}$. The pseudocode outlined in Algorithm 1 describes the main procedure to find invocable theorems. The process involves initializing a gym-like environment to interact with Lean and extracting the initial proof state for the candidate statement. Then, the algorithm iteratively tests whether one theorem can be used to rewrite or apply to the candidate theorem leveraging the instruction templates shown in Table 1. Suppose the feedback from the interactive environment is deemed valid according to predefined criteria, the algorithm adds the proof states before and after the tactic running together with the respective instruction to the set of invocable theorems $T_{inv}$. More information about this process is described in Appendix C.2.
@@
rw Equality invocable_theorem : $a=b$ or $a\iff b$
@@
apply Implication invocable_theorem : $a\implies b$
@@
Mutate statements. After obtaining the initial set of invocable theorems, we applied some filtering rules to $T_{inv}$ to improve the quality of the data and lower the complexity of mutating statements. With filtered invocable theorems, we construct new statements by replacing the components with their equivalent forms or antecedents. Since we use tactics in Lean to formulate the manipulations $\Phi$, most symbolic manipulations are bypassed to the Lean proof assistant. What remains is just parsing and replacing. Specifically, for the candidate statement $s$ and instruction $i$, we utilize its abstract syntax tree to pinpoint the exact location within the code that requires modification. Then we replace the corresponding parts with mutants parsing from the subsequent proof state generated by the execution of a specific tactic. The details of our algorithm are described in C.3.